OpenStudy (anonymous):

I will give a medal to whoever helps me. Solve. x/2-1/2=5

OpenStudy (jdoe0001):

multiply both sides by 2 \(\bf \cfrac{x}{2}-\cfrac{1}{2}=5\implies \cancel{2}\cdot \cfrac{x}{\cancel{2}}-\cancel{2}\cdot \cfrac{1}{\cancel{2}}=2\cdot 5\)

OpenStudy (anonymous):

Some of my answers are 4, 6, 11, and 16

OpenStudy (jdoe0001):

well, what did you get?

OpenStudy (jdoe0001):

dunno, see what you'd get from the simplication

OpenStudy (anonymous):

I don't get this could u help me plz

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=zNtqneXTSWM <---